# San Jose/Seattle, US https://bytedance.github.io/vidi-website/

### v1

@@ Abstract
Video has emerged as the primary medium for communication and creativity on the Internet, driving strong demand for scalable, high-quality video production. Vidi models continue to evolve toward next-generation video creation and have achieved state-of-the-art performance in multimodal temporal retrieval (TR). In its second release, Vidi2 advances video understanding with fine-grained spatio-temporal grounding (STG) and extends its capability to video question answering (Video QA), enabling comprehensive multimodal reasoning. Given a text query, Vidi2 can identify not only the corresponding timestamps but also the bounding boxes of target objects within the output time ranges. This end-to-end spatio-temporal grounding capability enables potential applications in complex editing scenarios, such as plot or character understanding, automatic multi-view switching, and intelligent, composition-aware reframing and cropping. To enable comprehensive evaluation of STG in practical settings, we introduce a new benchmark, VUE-STG, which offers four key improvements over existing STG datasets: 1) Video duration: spans from roughly $10$s to $30$ mins, enabling long-context reasoning; 2) Query format: queries are mostly converted into noun phrases while preserving sentence-level expressiveness; 3) Annotation quality: all ground-truth time ranges and bounding boxes are manually annotated with high accuracy; 4) Evaluation metric: a refined vIoU/tIoU/vIoU-Intersection scheme for multi-segment spatio-temporal evaluation. In addition, we upgrade the previous VUE-TR benchmark to VUE-TR-V2, achieving a more balanced video-length distribution and more user-style queries. Remarkably, the Vidi2 model substantially outperforms leading proprietary systems, such as Gemini 3 Pro (Preview) and GPT-5, on both VUE-TR-V2 and VUE-STG, while achieving competitive results with popular open-source models with similar scale on video QA benchmarks.
@@ 1 Introduction
As an inherently multimodal medium, video has become a dominant channel for online communication and information sharing. However, producing high-quality videos remains challenging for most users, particularly when performing trimming or editing operations on mobile devices. The first release of Vidi [team2025vidi] demonstrated strong temporal understanding across text, visual, and audio modalities, revealing its potential for automatic video trimming and understanding. To advance toward next-generation video creation systems capable of handling complex editing scenarios, Vidi models have been continually evolving to support more comprehensive multimodal perception and reasoning capabilities.
The second release, Vidi2, introduces for the first time an end-to-end spatio-temporal grounding (STG) capability, identifying not only the temporal segments corresponding to a text query but also the spatial locations of relevant objects within those frames. Notably, existing academic models [gu2024context, gu2025knowing, jin2022embracing, lin2023collaborative, su2021stvgbert, wasim2024videogrounding] and state-of-the-art industrial systems such as Gemini 3 Pro (Preview) [Gemini, comanici2025gemini] and GPT-5 [OpenAI2025GPT5SystemCard] are not yet capable of producing such fine-grained grounding results in a unified text output format (see Fig. 1). As illustrated in Fig. 2, the query “a man standing up from a kneeling position” represents a particularly difficult case involving multiple people in a dark scene. Vidi2 accurately localizes the corresponding time ranges and distinguishes the target person from others through precise bounding boxes in an end-to-end manner. It requires a comprehensive understanding of the visual and temporal dynamics of the described action. This fine-grained spatio-temporal perception highlights Vidi2’s potential to enable advanced editing workflows, such as plot-level understanding, automatic multi-view switching, and composition-aware reframing for professional video creation.
@@
To enable comprehensive evaluation of spatio-temporal grounding (STG) in realistic scenarios, we introduce a new benchmark, VUE-STG, featuring videos that range from $10$ seconds to $30$ minutes, substantially longer than existing STG datasets [chen2019weakly, tang2021human, yamaguchi2017spatio, zhang2020does]. Compared to prior work, VUE-STG includes videos and annotated tubes of much greater temporal extent, along with more complex, multimodal queries that often incorporate audio cues to assist temporal localization. All timestamps and bounding boxes are manually annotated with high precision, ensuring reliable spatio-temporal alignment. In addition, we upgrade the previous VUE-TR benchmark to VUE-TR-V2, featuring more human-authored, free-form queries and a more balanced video-length distribution with an increased proportion of long and ultra-long videos, making the task more realistic and challenging.
@@
On the VUE-STG benchmark, Vidi2 surpasses leading proprietary models (Gemini 3 Pro (Preview) and GPT-5) by a substantial margin, demonstrating state-of-the-art spatio-temporal grounding capability among large multimodal models. Compared with the previous version, Vidi2 also shows significant improvements in temporal retrieval, particularly for long videos. On the updated VUE-TR-V2 benchmark, it outperforms Gemini 3 Pro (Preview) [^2] Note that Gemini 3 Pro shows a notable improvement over Gemini-2.5-Pro-0325 on temporal retrieval. and GPT-5 by a wide margin across “Medium” to “Ultra-Long” video categories, confirming its leading performance in temporal reasoning. Furthermore, Vidi2 extends its capability to generic video question answering, marking a step toward a foundation model for comprehensive video understanding with strong spatio-temporal perception.
@@ 2 Model Overview
Vidi2 builds upon the foundation of Vidi [team2025vidi] with substantial upgrades in architecture, training data, and task capabilities, most notably the support for spatio-temporal grounding (STG) and video question answering (Video QA). These improvements collectively enable fine-grained multimodal understanding across text, visual, and audio modalities.
@@
Vidi2 retains the multimodal architecture of Vidi [team2025vidi], designed to jointly process text, visual, and audio inputs, while introducing key enhancements in both the encoder and LLM backbone (e.g., Gemma-3 [team2025gemma]). All results presented in this report are based on the 12B-parameter configuration. To achieve robust performance across videos of different lengths, we re-design the adaptive token compression strategy, improving the balance between short and long video representation efficiency. Additionally, a single image is treated as a $1$-second silent video, ensuring a unified encoding interface for both image and video inputs.
@@
Vidi2 follows a training pipeline similar to the previous Vidi version but scales up in both data volume and data diversity. While synthetic data remains essential for coverage and stability, increasing the proportion of real video data significantly enhances performance across all video-related tasks. The supervised fine-tuning (SFT) dataset for temporal retrieval has been expanded and refined in both quality and quantity. Furthermore, we incorporate additional generic QA data during SFT to strengthen open-ended reasoning on images and videos, improving the model’s versatility on downstream applications such as highlight detection, dense captioning, and video summarization.
@@
The previous Vidi [team2025vidi] model demonstrated the effectiveness of the temporal-aware multimodal alignment training stage (see Sec. 4 in [team2025vidi]) for temporal retrieval. Vidi2 further validates the generalization capability of this multimodal alignment strategy to generic video QA. By incorporating video QA data during the post-training phase, Vidi2 is able to answer both visual and auditory questions with high accuracy. We also observe that scaling up multimodal alignment data consistently improves overall video QA performance.
@@ 3 Evaluation Benchmark
To comprehensively evaluate the performance and generalization of Vidi2, we consider three representative categories of video understanding tasks: Spatio-Temporal Grounding (STG), Temporal Retrieval (TR), and Video Question Answering (Video QA). The first two are evaluated on our newly proposed benchmarks (i.e., VUE-STG and VUE-TR-V2), which are designed to assess fine-grained temporal and spatio-temporal reasoning under realistic editing and understanding scenarios. For video QA, we adopt widely used public benchmarks, including LVBench [wang2025lvbench], LongVideoBench [wu2024longvideobench], and VideoMME [fu2025video]. All of these video QA datasets employ a multiple-choice question format, which enables standardized, objective, and reproducible evaluation of multimodal reasoning performance. Together, these datasets provide a unified and comprehensive evaluation protocol that spans retrieval, grounding, and video reasoning.
@@ 3.1 VUE-STG
We introduce VUE-STG, a new Video Understanding Evaluation benchmark specifically designed to advance Spatio-Temporal Grounding in real-world scenarios. VUE-STG addresses four critical aspects often overlooked in prior academic benchmarks [chen2019weakly, tang2021human, yamaguchi2017spatio, zhang2020does]: video duration, query format, annotation quality, and metric design. To ensure the accurate performance signal, all ground-truth labels are manually curated, which yields significantly more accurate and consistent labels than existing benchmarks.
@@ 3.1.1 Data Distribution
Following VUE-TR [team2025vidi], VUE-STG is built using publicly available videos, and the annotations are made accessible for open research. As presented in Table 1, the benchmark consists of $1,600$ queries across $982$ videos, spanning over $204.79$ hours. It supports attribute-based slicing for fine-grained performance analysis as follows.
  • Video Duration. Unlike existing datasets limited to short videos, the video length varies from $10$ seconds to $30$ minutes, covering most of the duration ranges encountered in real-world scenarios. To enable duration-wise evaluation, we categorize videos into three balanced buckets following the definition of VUE-TR [team2025vidi]: ultra-short (<1<1 min), short (1101-10 mins), and medium (103010-30 mins).
@@ 3.2 VUE-TR-V2
Following the settings of VUE-TR [team2025vidi], we introduce an updated version, VUE-TR-V2, to provide a more balanced video length distribution and better query quality of the evaluation dataset.
@@
We follow the previous VUE-TR [team2025vidi] benchmark that uses the AUC (Area Under Curve) score of temporal-dimension precision, recall, and IoU (Intersection over Union) in the evaluation, denoted as $\bar{P}$, $\bar{R}$, $\bar{\text{IoU}}$ in Sec. 4. In particular, $\bar{\text{IoU}}$ is used as the primary metric.

### v2

@@ Abstract
Video has emerged as the primary medium for communication and creativity on the Internet, driving strong demand for scalable, high-quality video production. Vidi models continue to evolve toward next-generation video creation and have achieved state-of-the-art performance in multimodal temporal retrieval (TR). In its second release, Vidi2 advances video understanding with fine-grained spatio-temporal grounding (STG) and extends its capability to video question answering (Video QA), enabling comprehensive multimodal reasoning. Given a text query, Vidi2 can identify not only the corresponding timestamps but also the bounding boxes of target objects within the output time ranges. This end-to-end spatio-temporal grounding capability enables potential applications in complex editing scenarios, such as plot understanding, automatic view switching, etc. To enable comprehensive evaluation of STG in practical settings, we introduce a new benchmark, VUE-STG, which offers critical improvements over existing STG datasets in video duration, query format, annotation quality, and evaluation metric. In addition, we upgrade the previous VUE-TR benchmark to VUE-TR-V2, achieving a more balanced duration and query distribution. Remarkably, the Vidi2 model substantially outperforms leading proprietary systems, such as Gemini 3 Pro Preview and GPT-5, on both VUE-TR-V2 and VUE-STG, while achieving competitive results with popular open-source models with similar scale on video QA benchmarks. The latest Vidi2.5 offers significantly stronger STG capability and slightly better TR and Video QA performance over Vidi2 by RL training. This update also introduces a Vidi2.5-Think model to handle plot understanding with complex plot reasoning. To comprehensively evaluate the performance of plot understanding, we propose a new benchmark named VUE-PLOT with two tracks, Character and Reasoning. Notably, Vidi2.5-Think outperforms Gemini 3 Pro Preview on fine-grained character understanding with comparable performance on complex plot reasoning. Furthermore, we demonstrate the effectiveness of Vidi2.5 on a challenging real-world application, video editing planning. The post-trained variant Vidi-Edit can generate structured editing plans specifying narrative structure, audio attributes, and visual editing intent, illustrating its strong potential for complex real-world tasks.
@@ 1 Introduction
As an inherently multimodal medium, video has become a dominant channel for online communication and information sharing. However, producing high-quality videos remains challenging for most users, particularly when performing trimming or editing operations on mobile devices. The first release of Vidi [17] demonstrated strong temporal understanding across text, visual, and audio modalities, revealing its potential for automatic video trimming and understanding. To advance toward next-generation video creation systems capable of handling complex editing scenarios, Vidi models have been continually evolving to support more comprehensive multimodal perception and reasoning capabilities.
The second release, Vidi2 [18], introduces for the first time an end-to-end spatio-temporal grounding (STG) capability, identifying not only the temporal segments corresponding to a text query but also the spatial locations of relevant objects within those frames. Notably, existing academic models [4, 5, 9, 10, 13, 20] and state-of-the-art industrial systems such as Gemini 3 Pro Preview [7, 2] and GPT-5 [11] are not yet capable of producing such fine-grained grounding results in a unified text output format (see Fig. 1). As illustrated in Fig. 2, the query “a man standing up from a kneeling position” represents a particularly difficult case involving multiple people in a dark scene. Vidi2 accurately localizes the corresponding time ranges and distinguishes the target person from others through precise bounding boxes in an end-to-end manner. It requires a comprehensive understanding of the visual and temporal dynamics of the described action. This fine-grained spatio-temporal perception highlights Vidi2’s potential to enable advanced editing workflows, such as plot-level understanding, automatic multi-view switching, and composition-aware reframing for professional video creation.
@@
To enable comprehensive evaluation of spatio-temporal grounding (STG) in realistic scenarios, we introduce a new benchmark, VUE-STG, featuring videos that range from $10$ seconds to $30$ minutes, substantially longer than existing STG datasets [1, 14, 23, 24]. Compared to prior work, VUE-STG includes videos and annotated tubes of much greater temporal extent, along with more complex, multimodal queries that often incorporate audio cues to assist temporal localization. All timestamps and bounding boxes are manually annotated with high precision, ensuring reliable spatio-temporal alignment. In addition, we upgrade the previous VUE-TR benchmark to VUE-TR-V2, featuring more human-authored, free-form queries and a more balanced video-length distribution with an increased proportion of long and ultra-long videos, making the task more realistic and challenging.
@@
On the VUE-STG benchmark, Vidi2 surpasses leading proprietary models (Gemini 3 Pro Preview and GPT-5) by a substantial margin, demonstrating state-of-the-art spatio-temporal grounding capability among large multimodal models. Compared with the previous version, Vidi2 also shows significant improvements in temporal retrieval, particularly for long videos. On the updated VUE-TR-V2 benchmark, it outperforms Gemini 3 Pro PreviewNote that Gemini 3 Pro shows a notable improvement over Gemini-2.5-Pro-0325 on temporal retrieval. and GPT-5 by a wide margin across “Medium” to “Ultra-Long” video categories, confirming its leading performance in temporal reasoning. Furthermore, Vidi2 extends its capability to generic video question answering, marking a step toward a foundation model for comprehensive video understanding with strong spatio-temporal perception.
Figure 3: Example of plot understanding and reasoning tasks in VUE-PLOT. The Character track requires dense speaker localization and speech recognition, while the Reasoning track evaluates narrative understanding, character relationships, and editing techniques via multiple-choice visual question answering. Speakers’ faces and spoken lines are shown on the video for better visualization.
The latest Vidi2.5 introduces several major updates in Section 1, 3.3, 4.4, 5 to introduce new models, benchmarks and applications. Complementary results and description are added in Section 2, 4. The base Vidi2.5 model introduces multimodal Reinforcement Learning (RL) training with verifiable rewards and consistently boosts STG, TR, and Video QA performance. To further enhance multimodal reasoning capabilities for complex scenarios, e.g., plot understanding and reasoning, we introduce a thinking model for the Vidi2.5 model family. Inspired by DeepSeek-R1 [6] and OpenAI o1 [8], Vidi2.5-Think supports inference-time scaling to produce higher-quality responses. Unlike these models, Vidi2.5-Think is optimized using a combination of verifiable rewards and generative reward prediction, specifically designed to capture fine-grained details in videos and enable complex plot reasoning.
To comprehensively evaluate models on both fine-grained perception and high-level plot understanding, we introduce a new benchmark: VUE-PLOT. VUE-PLOT is formed by two tracks, Character and Reasoning. Character focuses on real-world videos and assesses fine-grained character perception by requiring models to densely localize the speaking individuals and recognize their spoken content. Reasoning targets high-level plot understanding and reasoning, challenging models to reason about narrative dynamics, character relationships, and editing techniques through multiple-choice visual question answering on informative video clips. Notably, our Vidi2.5-Think model outperforms Gemini 3 Pro on character grounding and speech recognition, while achieving competitive performance on complex plot reasoning tasks.
In addition to model and benchmark advances, Vidi2.5 is also applied to practical video creation workflows. We study video editing planning as a representative application, implemented via a post-trained variant, Vidi-Edit, which transforms raw video assets and optional user intents into structured editing plans. These plans explicitly specify narrative structure, voiceover content, audio attributes, and visual editing intent, serving as high-level specifications for downstream video creation. This application illustrates how Vidi2.5 can be used to support complex, real-world video editing workflows beyond standard video understanding settings.
@@ 2 Model Overview
Vidi2 builds on the foundation of Vidi [17] with substantial upgrades in architecture, training data, and task capabilities, most notably the support for spatio-temporal grounding (STG) and video question answering (Video QA). These improvements collectively enable fine-grained multimodal understanding across text, visual, and audio modalities.
@@
Vidi2 retains the multimodal architecture of Vidi [17], designed to jointly process text, visual, and audio inputs, while introducing key enhancements in both the encoder and LLM backbone (e.g., Gemma-3 [15]). All results presented in this report are based on the 12B-parameter configuration. To achieve robust performance across videos of different lengths, we re-design the adaptive token compression strategy, improving the balance between short and long video representation efficiency. Additionally, a single image is treated as a $1$-second silent video, ensuring a unified encoding interface for both image and video inputs.
@@
Vidi2 follows a training pipeline similar to the previous Vidi version but scales up in both data volume and data diversity. While synthetic data remains essential for coverage and stability, increasing the proportion of real video data significantly enhances performance across all video-related tasks. The supervised fine-tuning (SFT) dataset for temporal retrieval has been expanded and refined in both quality and quantity. Furthermore, we incorporate additional generic QA data during SFT to strengthen open-ended reasoning on images and videos, improving the model’s versatility on downstream applications such as highlight detection, dense captioning, and video summarization. Additional human annotated data is collected for RL training of Vidi2.5 on verifiable tasks.
@@
The previous Vidi [17] model demonstrated the effectiveness of the temporal-aware multimodal alignment training stage (see Section 4 in [17]) for temporal retrieval. Vidi2 further validates the generalizability of this multimodal alignment strategy to generic video QA. By incorporating video QA data during the post-training phase, Vidi2 is able to answer both visual and auditory questions with high accuracy. We also observe that scaling up multimodal alignment data consistently improves overall video QA performance.
Thinking Model:
Vidi2.5 introduces a new thinking model, Vidi2.5-Think, to enhance plot reasoning in videos via inference-time scaling. Vidi2.5-Think is trained using a highly customized reinforcement learning framework built on VERL [12]. The training data consist of diverse multimodal instruction-following examples covering video perception, audio understanding, narrative comprehension, and professional filming and editing techniques. These instructions include both multiple-choice visual question answering tasks and open-ended reasoning tasks. The model is optimized using a combination of verifiable rewards and large language models acting as judges to evaluate the quality of predicted rationales, generated responses, and their mutual consistency.
@@ 3 Evaluation Benchmark
To comprehensively evaluate the performance and generalization of Vidi2, we consider three representative categories of video understanding tasks: Spatio-Temporal Grounding (STG), Temporal Retrieval (TR), and Video Question Answering (Video QA). The first two are evaluated on our newly proposed benchmarks (i.e., VUE-STG and VUE-TR-V2), which are designed to assess fine-grained temporal and spatio-temporal reasoning under realistic editing and understanding scenarios. For video QA, we adopt widely used public benchmarks, including LVBench [19], LongVideoBench [21], and VideoMME [3]. All of these video QA datasets employ a multiple-choice question format, which enables standardized, objective, and reproducible evaluation of multimodal reasoning performance. Together, these datasets provide a unified and comprehensive evaluation protocol that spans retrieval, grounding, and video reasoning. We further introduce VUE-PLOT along with Vidi2.5 for complex plot understanding in real-world video editing scenarios.
@@ 3.1 VUE-STG
We introduce VUE-STG, a new Video Understanding Evaluation benchmark specifically designed to advance Spatio-Temporal Grounding in real-world scenarios. VUE-STG addresses four critical aspects often overlooked in prior academic benchmarks [1, 14, 23, 24]: video duration, query format, annotation quality, and metric design. To ensure the accurate performance signal, all ground-truth labels are manually curated, which yields significantly more accurate and consistent labels than existing benchmarks.
@@ 3.1.1 Data Distribution
Following VUE-TR [17], VUE-STG is built using publicly available videos, and the annotations are made accessible for open research. As presented in Table 1, the benchmark consists of $1,600$ queries across $982$ videos, spanning over $204.79$ hours. It supports attribute-based slicing for fine-grained performance analysis as follows.
  • Video Duration. Unlike existing datasets limited to short videos, the video length varies from $10$ seconds to $30$ minutes, covering most of the duration ranges encountered in real-world scenarios. To enable duration-wise evaluation, we categorize videos into three balanced buckets following the definition of VUE-TR [17]: ultra-short (<1<1 min), short (1101-10 mins), and medium (103010-30 mins).
@@ 3.2 VUE-TR-V2
Following the settings of VUE-TR [17], we introduce an updated version, VUE-TR-V2, to provide a more balanced video length distribution and better query quality of the evaluation dataset.
@@
We follow the previous VUE-TR [17] benchmark that uses the AUC (Area Under Curve) score of temporal-dimension precision, recall, and IoU (Intersection over Union) in the evaluation, denoted as $\bar{P}$, $\bar{R}$, $\bar{\text{IoU}}$ in Sec. 4. In particular, $\bar{\text{IoU}}$ is used as the primary metric.
3.3 VUE-PLOT
Duration Character Reasoning
# Videos # Segments # Bboxes # Videos # QAs
<<90s 34 547 1276 7 65
90-150s 248 5630 13422 56 521
150-210s 235 6517 15944 58 488
>>210s 29 860 2441 16 140
Total 546 13554 33083 137 1214
Table 5: Data distribution statistics for the VUE-PLOT benchmark. The benchmark covers a diverse range of video durations, from short clips (<90<90s) to longer sequences (>210>210s), with a total video duration of 22.45 hours for the Character track and 6.11 hours for the Reasoning track. This comprehensive coverage enables robust evaluation of models across various temporal scales. The Character track provides fine-grained grounding annotations (segments and bounding boxes), while the Reasoning track challenges models with complex reasoning tasks across multiple domains.
Figure 6: The speech duration distribution (left) and the bounding box area distribution (middle) of the Character track of the VUE-PLOT benchmark and task type distribution (right) of the Reasoning track.
The Character track of VUE-PLOT serves as the foundation for fine-grained video understanding, comprising $546$ videos with a total duration of $22.45$ hours. As detailed in Tab. 5, the videos are categorized into four duration intervals to ensure temporal diversity: <90<90s, 9090-$150$s, 150150-$210$s, and >210>210s. The distribution is well-balanced across the middle intervals, providing a solid basis for training and evaluation. Beyond video-level stats, the benchmark is densely annotated with $13,554$ speech segments and $33,083$ bounding boxes. The segment durations are predominantly short, as shown in Fig. 6 (left), requiring precise temporal localization. Meanwhile, the bounding boxes cover a range of object sizes, with $51.1\%$ Small ($5\%$ area), $41.3\%$ Medium ($5\%$-$20\%$ area), and $7.6\%$ Large ($20\%$ area) as shown in Fig. 6 (middle), challenging models to handle scale variation.
The reasoning track of the VUE-PLOT benchmark, designed for high-level reasoning, consists of $1,214$ QA pairs derived from a diverse subset of $137$ videos (totaling $6.11$ hours). The questions are distributed across five distinct task types, as shown in Fig. 6 (right): Narrative and Structural Understanding, Perception and Understanding, Professional Filming and Editing Techniques, Social Cognition and Knowledge Integration, Speech, Audio, and Sound Effect Reasoning. This diversity ensures that models are tested on various cognitive dimensions. The textual complexity is also significant, with an average question length of $24.8$ words, an average option length of $24.5$ words, necessitating strong multimodal reasoning capabilities.
Evaluation Metric:
We evaluate the performance of models on plot understanding and reasoning using different metrics. Temporal Grounding IoU ($\mathrm{tIoU}$) is used to evaluate the temporal alignment accuracy between predicted and ground-truth transcript segments, we match each ground-truth segment including the interested speaking characters to at most one predicted segment based on temporal Intersection-over-Union (IoU). The temporal grounding score is reported as the average IoU over all matched segment pairs:
$\mathrm{tIoU}=\frac{1}{N_{\mathrm{match}}}\sum_{i=1}^{N_{\mathrm{match}}}\mathrm{IoU}(s_{i}^{\mathrm{gt}},s_{i}^{\mathrm{pred}}).$
Word Error Rate ($\mathrm{WER}$) measures the speech transcription quality under correct temporal grounding. Specifically, we concatenate the text from all matched ground-truth segments and their corresponding predictions, and compute the Word Error Rate (WER) between the resulting corpora:
$\mathrm{WER}=\frac{S+D+I}{N},$
where $S$, $D$, and $I$ denote the numbers of substitutions, deletions, and insertions, respectively, and $N$ is the number of words in the ground-truth transcript.
Bounding Box IoU. For matched transcript segments that include spatial annotations, we evaluate visual grounding accuracy using bounding box IoU. Ground-truth and predicted boxes are temporally aligned based on minimal timestamp difference (i.e., 2020ms in our setting), and IoU is computed for each matched box pair. The final score is reported as the average bounding box IoU over all matched boxes:
$\mathrm{sIoU}=\frac{1}{N_{\mathrm{box}}}\sum_{j=1}^{N_{\mathrm{box}}}\mathrm{IoU}(b_{j}^{\mathrm{gt}},b_{j}^{\mathrm{pred}}).$
Multiple-Choice VQA Accuracy. To evaluate high-level plot understanding in a multiple-choice setting, we measure accuracy on multiple-choice visual question answering tasks. For each question, the model selects one answer from a fixed set of candidate options. The Multiple-Choice VQA Accuracy is defined as the proportion of questions for which the predicted answer exactly matches the ground-truth answer:
$\mathrm{Acc}_{\mathrm{MC}}=\frac{1}{N_{\mathrm{q}}}\sum_{i=1}^{N_{\mathrm{q}}}\mathbb{I}\!\left(a_{i}^{\mathrm{pred}}=a_{i}^{\mathrm{gt}}\right),$
where $N_{\mathrm{q}}$ denotes the total number of multiple-choice questions, $a_{i}^{\mathrm{pred}}$ and $a_{i}^{\mathrm{gt}}$ represent the predicted and ground-truth answers for the $i$-th question, and $\mathbb{I}(\cdot)$ is the indicator function.